OpenStudy (anonymous):

tangent plane of (x^2/a^2)+ (y^2/b^2)+(z^2/c^2)=1 ?

OpenStudy (anonymous):

\[xx ^{*} /a ^{2} + yy^*/b^2 + zz^*/c^2 =1 ????\]

OpenStudy (anonymous):

tangent plane at what point?

OpenStudy (anonymous):

at a general point x^* y^* z^*

OpenStudy (anonymous):

okay change of answer it is x^*/a^2 +y^*/b^2 +z^*/c^2 =1 i think thats the correc tangent plane eq

OpenStudy (anonymous):

you need to take derivative of the parabola to get the tangent

OpenStudy (anonymous):

a parabola? btw its cal 2 and vector algebra

OpenStudy (anonymous):

x y and z are independent of each other.

OpenStudy (anonymous):

scrap that last post.

OpenStudy (anonymous):

rsaad, you there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

You can find the equation for the tangent plane at the point (x_0,y_0,z_0) by taking the gradient of your function, evaluating at that point (the grad. will give you a vector that is perpendicular to the plane) and sub. in those values into the equation for a plane.

OpenStudy (anonymous):

So...

OpenStudy (anonymous):

i did calcu;at the equation of tangent plane.

OpenStudy (anonymous):

\[f(x,y,z)=\frac{x^2}{a^2}+\frac{y^2}{b^2}+\frac{z^2}{c^2}-1=0\]

OpenStudy (anonymous):

Oh...what's wrong then?

OpenStudy (anonymous):

the problem i s i am getting the volume in negative.. =(

OpenStudy (anonymous):

\[\nabla f=(\frac{2x}{a^2},\frac{2y}{b^2},\frac{2z}{c^2})\]

OpenStudy (anonymous):

Volume? You don't need volume here.

OpenStudy (anonymous):

i took the triple product. should i attach my cal. ?

OpenStudy (anonymous):

Wait and see what I get, then you can compare.

OpenStudy (anonymous):

i need a volume of a tetrahedron which is formed when x=0,y=0,z=0 and with thi stangent planbe.

OpenStudy (anonymous):

At the point (x_0,y_0,z_0), the normal to the tangent plane will be\[\nabla f(x_0,y_0,z_0)=(\frac{2x_0}{a^2},\frac{2y_0}{b^2},\frac{2z_0}{c^2})\]and so, at the same point, the equation of the tangent plane will be,\[\frac{2x_0}{a^2}(x-x_0)+\frac{2y_0}{b^2}(y-y_0)+\frac{2z_0}{c^2}(z-z_0)=0\]that is,\[\frac{x_0}{a^2}(x-x_0)+\frac{y_0}{b^2}(y-y_0)+\frac{z_0}{c^2}(z-z_0)=0\]

OpenStudy (anonymous):

i got these coordinates \[(-a ^{2}/x , b ^{2}/y, 0), (-a ^{2}/x,0,c ^{2}/z),(-a ^{2}/x,0,0)\]

OpenStudy (anonymous):

yes and on simplifying that equation you get:\[xx _{o}^{}/a^2 +yy _{o}^{}/b^2 +zz _{o}^{}/c^2 =1\]

OpenStudy (anonymous):

then i used the tripple product to get vol. but its negative

OpenStudy (anonymous):

Yeah, see, I'm thinking the plane has to form one side of the tetrahedron. Given that, it will cut the x-, y- and z-axes at the points,\[(\frac{a^2}{x_0},0,0)\]\[(0,\frac{b^2}{y_0},0)\]\[(0,0,\frac{c^2}{z_0})\]

OpenStudy (anonymous):

hey why did you take -x,-y,-z? how did you know its negative?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

as in i took x y z . that is the positive axis. why did u take the negative ones?

OpenStudy (anonymous):

I haven't taken negative axes.

OpenStudy (anonymous):

sorry my bad i was jumping to the triple product... so with the coordinates u ve mentioned, i should be taking triple product of these? or as i have done in my calculaatios in the attachment?

OpenStudy (anonymous):

Well, it's hard for me to say because I don't have the entire question. I'm only piecing together the fact you have a tangent plane that's to form one surface of a tetrahedron, and with the other three sides lying in the x-y, x-z and x-y planes.

OpenStudy (anonymous):

The triple product is only going to find the volume of a parallelopiped.

OpenStudy (anonymous):

ok its find the min vol bounded by the planes, xy, yz,xz, and a plane tangent to the folowing: x^2/a^2+(y/b)^2 + (z/c)^2 =1

OpenStudy (anonymous):

Okay, I guessed right.

OpenStudy (anonymous):

whats wrong with my calculations =(

OpenStudy (anonymous):

I'd have to do the problem to see that. When do you need this? I have some things I need to take care of.

OpenStudy (anonymous):

the attachment is not opening here.

OpenStudy (anonymous):

it finally opened.

OpenStudy (anonymous):

its hw question. anyway i'll further brood on it.

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

ok. if i get a chance, i'll come back to it :)

OpenStudy (anonymous):

rsaad, if you're there, I have the answer for you.

OpenStudy (anonymous):

It's a long problem...

OpenStudy (anonymous):

There are a couple of ways of doing this (e.g. Lagrange multipliers), but substitution works fine too.

OpenStudy (anonymous):

I'll upload the first sheet where I've found the appropriate determinant using the coordinates I found before:

OpenStudy (anonymous):

Now, the x, y and z are points that satisfy the equation of your surface, so we may eliminate z as\[z=c \sqrt{1-\frac{x^2}{a^2}+\frac{y^2}{b^2}}\]

OpenStudy (anonymous):

and sub. this in for z in the volume equation. Now we have a function in x and y only. We can find the extrema by finding\[\frac{\partial V}{\partial x}=0\]and\[\frac{\partial V}{\partial y}=0\]

OpenStudy (anonymous):

lol, loki, you've got a neat writing

OpenStudy (anonymous):

You may solve for x^2 in the first to find\[x^2=\frac{(b^2-y^2)a^2}{2b^2}\]

OpenStudy (anonymous):

hehe, thanks :)

OpenStudy (anonymous):

^_^ np

OpenStudy (anonymous):

and sub. this into the partial derivative found for y to find:\[y^2=\frac{b^2}{3}\]Substituting this back into the first expression for x^2 we have,\[x^2=\frac{a^2}{3}\]

OpenStudy (anonymous):

To find z, we only need to realise that z is constrained by the surface, so substituting these in, we find,\[z^2=c^2(1-\frac{x^2}{a^2}-\frac{y^2}{b^2})=c^2(1-\frac{a^2/3}{a^2}-\frac{b^2/3}{b^2})=\frac{c^2}{3}\]

OpenStudy (anonymous):

Substituting these into the expression for volume, we have\[V=\frac{1}{6}\frac{a^2b^2c^2}{xyz}=\frac{1}{6}\frac{a^2b^2c^2}{\frac{a}{\sqrt{3}}\frac{b}{\sqrt{3}}\frac{c}{\sqrt{3}}}=\frac{\sqrt{3}}{2}abc\]

OpenStudy (anonymous):

You have to check that these values for x and y do yield a maximum in V by using the appropriate test\[f_{xx}<0\]and\[f_{xx}f_{yy}-f_{xy}^2>0\]

OpenStudy (anonymous):

And, um, give us a fan point if you haven't already ;p

OpenStudy (anonymous):

thanks lokisan =) but i've a question.why did you use the surface equation when actually it is the tangent plane which is bounding the tetrahedron?

OpenStudy (anonymous):

surface eq to eliminate z. why not the tangent plane equation? and also, why in calculating the detrminent, u used a-b, b-c,c-d where d was 0,0,0. on one website, that i consukted the formula was d-a,d-b,d-c... i am confused.

OpenStudy (anonymous):

You have to think of the surface equation as being the set of values x, y and z are allowed to take. Remember, the x, y, and z used in the volume formula were obtained from the plane equation, and that particular point was the point on the surface where the tangent plane was generated.

OpenStudy (anonymous):

I dropped the notation,\[V=\frac{1}{6}\frac{a^2b^2c^2}{x_0y_0z_0}\]for\[V=\frac{1}{6}\frac{a^2b^2c^2}{xyz}\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

You're not convinced...

OpenStudy (anonymous):

what about my 2nd q:and also, why in calculating the detrminent, u used a-b, b-c,c-d where d was 0,0,0. on one website, that i consukted the formula was d-a,d-b,d-c... i am confused.

OpenStudy (anonymous):

hang on...brb

OpenStudy (anonymous):

well yes. cuz when i plotted the tetrahedron, the sid was flanked by tangent PLANE. so why not set of points satisfied by that tangent plane? i mean if we look at the surface as a whole, wont that be misleading?

OpenStudy (anonymous):

For the determinant, I used the formula (1/6)·det(a−b, b−c, c−d), where a, b, c and d are the vertices. This one makes no specific mention of assuming a vertex (e.g. d) to be the origin.

OpenStudy (anonymous):

but like i said, i checked internet earlier and there it was mentioned d=000

OpenStudy (anonymous):

Hang on, this is easy to deal with. I want to talk about your other question.

OpenStudy (anonymous):

ok just tell me if we have found the 3 vertices, we simply subtract each of them to get 3 eqs?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

You asked, "so why not set of points satisfied by that tangent plane? i mean if we look at the surface as a whole, wont that be misleading?"

OpenStudy (anonymous):

Okay...the volume is determined by the tangent plane, and the tangent plane is determined by the surface. You're being asked to find the specific coordinates (i.e. point on surface which generates the tangent plane) that leads to a tangent plane that gives you the maximal volume. That's why the x, y and z must satisfy the surface equation, because we're generating tangent planes for various points (x,y,z) on the surface. We can't choose arbitrary x, y AND z, since these won't necessarily be on the surface, and therefore, won't give us a tangent plane to the surface. We can choose x and y, but the surface equation will dictate the allowable z we may take.

OpenStudy (anonymous):

O......!!! now i see! =) thanks!!!

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

Now, as for the equation for the volume...I used a formula from one of my books, and I've also seen it on wikipedia: http://en.wikipedia.org/wiki/Tetrahedron Look under 'volume'.

OpenStudy (anonymous):

You can derive this for yourself if you need convincing by understanding that the volume of any pyramid is 1/3 x base x perpendicular height.

OpenStudy (anonymous):

ahan. let me see

OpenStudy (anonymous):

right. so then it follows what you have done / thanks a lot! btw which uni r u from? i m curious 'cuz u r able to answer all my questions!

OpenStudy (anonymous):

A very old university... ;)

OpenStudy (anonymous):

IV league?

OpenStudy (anonymous):

Were all your questions answered?

OpenStudy (anonymous):

well yes. i'd posted some problems earlier and u solved them and so i'd fan-ed you already =)

OpenStudy (anonymous):

Ah...good...you caught me just in time, too...I was about to log out -.- Happy mathing :D

OpenStudy (anonymous):

=D lucky me! hey come on which uni r u from? i m curious! caltech?

OpenStudy (anonymous):

no... :)

OpenStudy (anonymous):

then? iv league?

OpenStudy (anonymous):

I'm a citizen of the UK

OpenStudy (anonymous):

oxford? cambridge?

OpenStudy (anonymous):

Well, it's possibly one of the two ;)

OpenStudy (anonymous):

aha! no doubt you are really good at math! =P ok thanks a lot!

OpenStudy (anonymous):

no probs :D